I have 3 questions dealing with linear equations, slopes, etc... that I can not figure out.
Can someone show me the steps on how to solve these problems, please???
1. Find the slope of y=-8x+4 ( I came up with -8, but not sure if that's correct)
2. Find the intercepts for the graph of the given equation x=-6y
3. Find an equation of the line containing the given pair of points (-2,-7) and (-9, -8)

Answer by stanbon(75887)

1. Find the slope of y=-8x+4 ( I came up with -8, but not sure if that's correct)
----
You are correct.
-----------------------------------
2. Find the intercepts for the graph of the given equation x=-6y
x-intercept occurs when y = 0.
x = -6*0
x = 0
So, x-intercept is (0,0)
---
Y-intercept occurs when x = 0.
y-intercept is (0,0)
=====================================
3. Find an equation of the line containing the given pair of points (-2,-7) and (-9, -8)
---
slope = (-7--8)/(-2--9) = 1/7
---
Form: y = mx + b
Solve for "b":
-7 = (1/7)(-2) + b
-7 = -2/7 + b
b = -49/7+2/7
b = -47/7
----
Equation:
y = (1/7)x - (47/7)
========================
Cheers,
Stan H.